OpenStudy (anonymous):

Consider the functions f(x) = 2x2 + 1 and g(x) = x + 2. What is the value of f[g(3)]?

OpenStudy (amistre64):

whats the value of g(3)?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

yes is not a number. what is the value of g(3)?

OpenStudy (amistre64):

since: g(x) = x + 2 what is the value of g(3) ?

OpenStudy (anonymous):

i dont know thats why im asking

OpenStudy (amistre64):

the notation means that when x=3, what does the value of the function of g take on? x+2 = ? , when x = 3

OpenStudy (anonymous):

5 but thats not one of the answers it gives me these choices 101, 51, 21, 11

OpenStudy (amistre64):

we havent gotten to an answer yet, that will be in the last step. The process to get to that answer is what we are working thru at the moment.

OpenStudy (amistre64):

so we have discovered the g(3) = 5, we can use this to our advantage f(g(3)) then means f(5)

OpenStudy (amistre64):

f(x) = 2x2 + 1 f(g(3)) = f(5) f(5) means what is the value of the function of f when its x is equal to 5 2x2 + 1 = ? when x = 5?

OpenStudy (amistre64):

feel free to ask questions if this doesnt make sense.

OpenStudy (anonymous):

51 (:

OpenStudy (amistre64):

51 is good :)

OpenStudy (anonymous):

thank you